# Linear Transformers are Versatile In-Context Learners

### v1

@@ 5.1 ${}^{++}$ and least squares solver
Given $(x_{1},y_{1}),...,(x_{n},y_{n}),(x_{t},0)$ where $\Sigma$ has eigenvalues in the range $[\nu,\mu]$ with a condition number $\kappa=\nu/\mu$. Let $w^{*}$ be the optimal solution to least squares problem (4), then there exists hyperparameters for ${}^{++}$ algorithm that outputs $\hat{y}$ with accuracy $|\hat{y}-\langle x_{t},w^{*}\rangle|\leq\epsilon\|x_{t}\|\|w^{*}\|$ in $l=O(\log\kappa+\log\log 1/\epsilon)$ steps. In particular that implies there exists an $l$-layer linear transformer that can solve this task.
@@
Figure 1: In-context learning performance for noisy linear regression problem across models with different number of layers and $\sigma_{max}$ for $\sigma_{\tau}\sim U(0,\sigma_{max})$. Each marker corresponds to a separately trained model with a given number of layers. Models with diagonal attention weights (Diag) match those with full attention weights (Full). Models specialized on a fixed noise (${}^{++}$) perform poorly, similar to a Ridge Regression solution with a constant noise (ConstRR). Among the baselines, only tuned exact Ridge Regression solution (TunedRR) is comparable with linear transformers.
5.2 Understanding $w_{yy}$: adaptive rescaling
If a layer only has $w^{l}_{yy}\neq 0$, it has a rescaling effect. The amount of scaling is related to the amount of noise added in a model selection setting. The update rule for this layer is:
$y^{l+1}_{i}=\left(1+w^{l}_{yy}\lambda^{l}\right)y^{l}_{i}.$
As we can see, this rescales every $y$ by a factor that depends on $\lambda^{l}$. When $w^{l}_{yy}<0$, this results in a shrinking of the output based on the norm of $y$ in the previous layer. This can be helpful for the mixed noise variance problem, because the ridge regression solution scales the least squares solution by a factor that depends on the noise level.
Specifically, assuming $\Sigma\approx\mathbb{E}[\Sigma]=nI$, the ridge regression solution becomes $w^{*}_{\sigma^{2}}\approx\frac{n}{n+\sigma^{2}}w^{*}$, which is exactly a scaled version of the OLS solution. Further, when noise is larger, the scaled factor is smaller, which agrees with the behavior of a negative $w_{yy}$.
5.3 Understanding $w_{xy}$: adapting step-sizes
The final term in the diagonal model, $w_{xy}$, has a more complicated effect. Since it changes only the $x$-coordinates, it does not have an immediate effect on $y$. To understand how it influences the $y$ we consider a simplified two-step process, where the first step only has $w_{xy}\neq 0$ and the second step only has $w_{yx}\neq 0$ (so the second step is just doing one step of gradient descent). In this case, the first layer will update the $x_{i}$’s as:
$x_{i}^{1}$ $=x_{i}+y_{i}w_{xy}{\sum_{j=1}^{n}}y_{j}x_{j}$
$=x_{i}+w_{xy}y_{i}{\sum_{j=1}^{n}}(\langle w^{*},x_{j}\rangle+r_{j})x_{j}$
$=x_{i}+w_{xy}y_{i}\Sigma w^{*}$
$=x_{i}+w_{xy}(\langle w^{*},x_{i}\rangle+r_{i})\Sigma w^{*}$
$=(I+w_{xy}\Sigma w^{*}(w^{*})^{\top})x_{i}+w_{xy}r_{i}\Sigma w^{*}.$
There are two effects of the $w_{xy}$ term, one is a multiplicative effect on $x_{i}$, and the other is an additive term that makes $x$-output related to the residual $r_{i}$. The multiplicative step in $x_{i}$ has an unknown preconditioning effect. For simplicity we assume the multiplicative term is small, that is:
$x_{i}^{1}\approx x_{i}+w_{xy}r_{i}\Sigma w^{*};\quad x_{t}^{1}\approx x_{t}.$
@@ 5.3 Understanding $w_{xy}$: adapting step-sizes
$y_{t}^{2}$ $=y_{t}+w_{yx}\sum_{i=1}^{n}y_{i}(x_{i}^{1})^{\top}x_{t}$
$\approx y_{t}+w_{yx}[\sum_{i=1}^{n}y_{i}x_{i}+w_{xy}\sum_{i=1}^{n}y_{i}r_{i}\Sigma w^{*}]x_{t}$
$=y_{t}+w_{yx}(1+w_{xy}\sum_{i=1}^{n}r_{i}^{2})(\Sigma w^{*})^{\top}x_{t}.$
Here we used the properties of residual $r_{i}$ (in particular $\sum_{i}y_{i}x_{i}=\Sigma w^{*}$, and $\sum_{i}y_{i}r_{i}=\sum_{i}r_{i}^{2}$). Note that $(\Sigma w^{*})^{\top}x_{t}$ is basically what a gradient descent step on the original input should do. Therefore effectively, the two-layer network is doing gradient descent, but the step size is the product of $-w_{yx}$ and $(1+w_{xy}\sum_{i}r_{i}^{2})$. The factor $(1+w_{xy}\sum_{i}r_{i}^{2})$ depends on the level of noise, and when $w_{xy},w_{yx}<0$, the effective step size is smaller when there is more noise. This is especially helpful in the model selection problem, because intuitively one would like to perform early-stopping (small step sizes) when the noise is high.
Figure 3: Per-variance profile of models behavior for uniform noise variance $\sigma_{\tau}\sim U(0,\sigma_{max})$. Top two rows: 7-layer models with varying $\sigma_{max}$. Bottom row: models with varying numbers of layers, fixed $\sigma_{max}=5$. In-distribution noise is shaded grey.
@@ 6 Experiments
  • ${}^{++}$. An even more restricted diagonal variant defined in (11).
For each experiment, we train each linear transformer modifications with a varying number of layers ($1$ to $7$) using using Adam optimizer for 200 000200000200\,000200 000 iterations with a learning rate of $0.0001$ and a batch size of 2 04820482\,0482 048. In some cases, especially for large number of layers, we had to adjust learning rate to prevent stability issues. We report the best result out of $5$ runs with different training seeds. We used $N=20$ in-context examples in $D=10$ dimensions. We evaluated the algorithm using 100 000100000100\,000100 000 novel sequences.
@@
Figure 4: In-context learning performance for noisy linear regression problem across models with different number of layers for conditional noise variance $\sigma_{\tau}\in\{1,3\}$ and $\sigma_{\tau}\in\{1,3,5\}$. Top row: loss for models with various number of layers and per-variance profile for models with 7 layers. Bottom row: Per-variance profile of the model across different numbers of layers. In-distribution noise is shaded grey.
@@
We begin with the uniform noise variance. Fig. 1 shows the performance of different models trained with varying numbers of layers and noise thresholds $\sigma_{max}$. Notably, Full and Diag achieve comparable performance across different numbers of layers and different $\sigma_{max}$. On the other hand, ${}^{++}$ converges to a higher value, closely approaching the performance of the ConstRR baseline.
@@
We tested an additional hypothesis that linear transformers trained on a mixed noise variance problem result in an iterative algorithm, with each layer corresponding to one iteration with prediction $y^{l}_{n+1}$. While we minimize the loss only for the final layer’s prediction $y^{L}_{n+1}$, we also track the loss for each intermediate prediction $y^{l}_{n+1}$. Fig. 2 illustrates these intermediate iterations for models with different numbers of layers, trained on a uniform mixed noise variance problem with $\sigma_{max}=5$.
We observe that ${}^{++}$ gradually decreases the loss after each layer after the second (surprisingly, the first layer does not improve predictions in models with more than 2 layers). Diag and Full behave more irregularly, with dramatic loss decrease happening in the first and final layers. Intermediate layers improve the prediction, but not consistently. This behavior suggests that these models might be accumulating information in the $x$ components of the prediction before making the prediction at the last layer.
The top of Fig. 3 offers a detailed perspective on performance of 7-layer models and the baselines. Here, we computed per-variance profiles across noise variance range from 0 to $\sigma_{max}+1$. We can see that poor performance of ${}^{++}$ comes from its inability to estimate well across the full noise variance range. Its performance closely mirrors to ConstRR, suggesting that ${}^{++}$ under the hood might also be estimating a single constant variance for all the data.
AdaRR perfectly estimates problems with no noise, but struggles more as noise variance grows. TunedRR improves the estimation a little by incorporating $\sigma_{max}$ into its tunable parameters. However, its prediction suffers in the middle range. Full and Diag are very closely comparable and perform well across the entire spectrum of noise variance. While more research is needed to definitively confirm or deny their equivalence, we believe that these models are actually not identical despite their similar performance.
At the bottom of Fig. 3 we fixed the noise variance to $\sigma_{max}=5$ and show a per-variance profile for models with different layers. 2-layer models for Full and Diag behave similarly to ${}^{++}$, modeling only a single noise variance in the middle. However, the results quickly improve across the entire noise spectrum for 3 or more layers. In contrast, ${}^{++}$ quickly converges to a suboptimal solution.
@@
Fig. 4 shows a notable difference between Diag and Full models for categorical noise variance $\sigma_{\tau}\in\{1,3\}$. This could stem from a bad local minima, or suggest a fundamental difference between the models for this problem. Interestingly, from per-variance profiling we see that Diag extrapolates better for variances not used for training, while Full, despite its lower in-distribution error, performs worse on unseen variances.
@@
For easy comparison, we also summarize the results of all methods and baselines in Table 1 in the Appendix.
@@ 7 Conclusions
We show that even linear transformers have remarkable capacity to handle challenging in-context learning problems. We prove that any linear transformer maintains an implicit linear model of the input data. Under certain restrictions, this model can be interpreted as a complex variant of preconditioned gradient descent with momentum-like behavior.
When trained on noisy linear regression problems with unknown noise variance, linear transformers surpass standard baselines and uncover a sophisticated optimization algorithm, incorporating noise-aware step-size adjustments and rescaling based on noise levels.
Our findings underscore the surprising ability of linear transformers to discover novel optimization algorithms when exposed to the right problems. This opens up exciting possibilities for future research, such as automated algorithm discovery using transformers or possible generalization to other problems. Beyond linear regression, we hope that our work will inspire further exploration into the ability of transformers to learn and represent optimization algorithms.
Our work highlights that even seemingly simple models, such as linear transformers, can embody remarkable complexity in the optimization algorithms they implicitly learn. While more work is needed, we hope that our paper can contribute to understanding the mechanisms behind in-context learning.
8 Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Omitted proofs from Section 4
@@
Next we will prove Lemma 4.3. This lemma is in fact a corollary of Lemma 4.2. We first give a more detailed version which explicitly state the unknown matrices $\Lambda^{l},\Gamma^{l},\Pi^{l},\Phi^{l}$:
@@
$u^{l+1}$ $=\left((1+w^{l}_{xy}(a^{l})^{2}\rho)I+w^{l}_{xx}\Sigma^{l}\right)u^{l}$
$+a^{l}w^{l}_{xy}\left(M^{l}+a^{l}u^{l}(w^{*})^{\top}\right)\Sigma\left(a^{l}w^{*}-w^{l}\right),$
$w^{l+1}$ $=(1+w^{l}_{yy}\lambda^{l})w^{l}$
$-w^{l}_{yx}(M^{l})^{\top}(M^{l}+a^{l}u^{l}(w^{*})^{\top})\Sigma(a^{l}w^{*}-w^{l})$
$-a^{l}\rho w^{l}_{yx}(M^{l})^{\top}u^{l},$
@@
$=\left(\begin{array}[]{cc}w^{l}_{xx}\Sigma^{l}&w^{l}_{xy}\alpha^{l}\\ w^{l}_{yx}(\alpha^{l})^{\top}&w^{l}_{yy}\lambda^{l}\end{array}\right).$
This implies that $A^{l}=w^{l}_{xx}\Sigma^{l}$, $b^{l}=w^{l}_{xy}\alpha^{l}$, $c^{l}=w^{l}_{yx}\alpha^{l}$ and $d^{l}=w^{l}_{yy}\lambda^{l}$. Next we rewrite $\alpha^{l}$:
@@
Now Lemma A.1 implies Lemma 4.3 immediately by setting $\Lambda^{l}=-w^{l}_{xy}(a^{l})^{2}\rho)I-w^{l}_{xx}\Sigma^{l}$, $\Gamma^{l}=a^{l}w^{l}_{xy}\left(M^{l}+a^{l}u^{l}(w^{*})^{\top}\right)$, $s^{l}=tu{w}_{yy}\lambda^{l}$, $\Pi^{l}=w^{l}_{yx}(M^{l})^{\top}(M^{l}+a^{l}u^{l}(w^{*})^{\top})$ and $\Phi^{l}=a^{l}\rho w^{l}_{yx}(M^{l})^{\top}$.
Appendix B Omitted proof for Theorem 5.1
In this section we prove Theorem 5.1 by finding hyperparameters for ${}^{++}$ algorithm that solves least squares problems with very high accuracy. The first steps in the construction iteratively makes the data $x_{i}$’s better conditioned, and the last step is a single step of gradient descent. The proof is based on several lemma, first we observe that if the data is very well-conditioned, then one-step gradient descent solves the problem accurately:
Lemma B.1.
@@
Next we show that by applying just the preconditioning step of ${}^{++}$, one can get a well-conditioned $x$ matrix very quickly. Note that the $\Sigma$ matrix is updated as $\Sigma\leftarrow(I-\gamma\Sigma)\Sigma(I-\gamma\Sigma)$, so an eigenvalue of $\lambda$ in the original $\Sigma$ matrix would become $\lambda(1-\gamma\lambda)^{2}$. The following lemma shows that this transformation is effective in shrinking the condition number
Lemma B.2.

### v2

@@ 5.1 GD++ and least squares solver
Given $(x_{1},y_{1}),...,(x_{n},y_{n}),(x_{t},0)$ where $\Sigma$ has eigenvalues in the range $[\nu,\mu]$ with a condition number $\kappa=\nu/\mu$. Let $w^{*}$ be the optimal solution to least squares problem (4), then there exists hyperparameters for GD++ algorithm that outputs $\hat{y}$ with accuracy $|\hat{y}-\langle x_{t},w^{*}\rangle|\leq\epsilon\|x_{t}\|\|w^{*}\|$ in $l=O(\log\kappa+\log\log 1/\epsilon)$ steps. In particular that implies there exists an $l$-layer linear transformer that can solve this task.
@@
5.2 Understanding $\omega_{yy}$: adaptive rescaling
If a layer only has $\omega^{l}_{yy}\neq 0$, it has a rescaling effect. The amount of scaling is related to the amount of noise added in a model selection setting. The update rule for this layer is:
$y^{l+1}_{i}=\left(1+\omega^{l}_{yy}\lambda^{l}\right)y^{l}_{i}.$
This rescales every $y$ by a factor that depends on $\lambda^{l}$. When $\omega^{l}_{yy}<0$, this shrinks of the output based on the norm of $y$ in the previous layer. This is useful for the mixed noise variance problem, as ridge regression solution scales the least squares solution by a factor that depends on the noise level.
Specifically, assuming $\Sigma\approx\mathbb{E}[\Sigma]=nI$, the ridge regression solution becomes $w^{*}_{\sigma^{2}}\approx\frac{n}{n+\sigma^{2}}w^{*}$, which is exactly a scaled version of the OLS solution. Further, when noise is larger, the scaled factor is smaller, which agrees with the behavior of a negative $\omega_{yy}$.
We can show that using adaptive scaling $\omega_{yy}$ even a 2-layer linear transformer can be enough to solve a simple example of categorical mixed noise variance problem $\sigma_{\tau}\in\{\sigma_{1},\sigma_{2}\}$ and $n\rightarrow\infty$:
Theorem 5.2.
Suppose the input to the transformer is $(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n}),(x_{q},0)$, where $x_{i}\sim N(0,\frac{1}{n}I)$, $y_{i}=w^{\top}x_{i}+\xi_{i}$. Here $\xi_{i}\sim N(0,\sigma^{2})$ is the noise whose noise level $\sigma$ can take one of two values: $\sigma_{1}$ or $\sigma_{2}$. Then as $n$ goes to +$\infty$+ ∞, there exists a set of parameters for two-layer linear transformers such that the implicit $w^{2}$ of the linear transformer converges to the optimal ridge regression results (and the output of the linear transformer is $-\langle w^{2},x_{q}\rangle$). Further, the first layer only has $\omega_{yx}$ being nonzero and the second layer only has $\omega_{yy}$ being nonzero.
5.3 Understanding $\omega_{xy}$: adapting step-sizes
The final term in the diagonal model, $\omega_{xy}$, has a more complicated effect. Since it changes only the $x$-coordinates, it does not have an immediate effect on $y$. To understand how it influences the $y$ we consider a simplified two-step process, where the first step only has $\omega_{xy}\neq 0$ and the second step only has $\omega_{yx}\neq 0$ (so the second step is just doing one step of gradient descent). In this case, the first layer will update the $x_{i}$’s as:
$x_{i}^{1}$ $=x_{i}+y_{i}\omega_{xy}{\sum_{j=1}^{n}}y_{j}x_{j}$
$=x_{i}+\omega_{xy}y_{i}{\sum_{j=1}^{n}}(\langle w^{*},x_{j}\rangle+r_{j})x_{j}$
$=x_{i}+\omega_{xy}y_{i}\Sigma w^{*}$
$=x_{i}+\omega_{xy}(\langle w^{*},x_{i}\rangle+r_{i})\Sigma w^{*}$
$=(I+\omega_{xy}\Sigma w^{*}(w^{*})^{\top})x_{i}+\omega_{xy}r_{i}\Sigma w^{*}.$
There are two effects of the $\omega_{xy}$ term, one is a multiplicative effect on $x_{i}$, and the other is an additive term that makes $x$-output related to the residual $r_{i}$. The multiplicative step in $x_{i}$ has an unknown preconditioning effect. For simplicity we assume the multiplicative term is small, that is:
$x_{i}^{1}\approx x_{i}+\omega_{xy}r_{i}\Sigma w^{*};\quad x_{t}^{1}\approx x_{t}.$
@@ 5.3 Understanding $\omega_{xy}$: adapting step-sizes
Figure 1: In-context learning performance for noisy linear regression problem across models with different number of layers and $\sigma_{max}$ for $\sigma_{\tau}\sim U(0,\sigma_{max})$. Each marker corresponds to a separately trained model with a given number of layers. Models with diagonal attention weights (Diag) match those with full attention weights (Full). Models specialized on a fixed noise (GD++) perform poorly, similar to a Ridge Regression solution with a constant noise (ConstRR). Among the baselines, only tuned exact Ridge Regression solution (TunedRR) is comparable with linear transformers.
$y_{t}^{2}$ $=y_{t}+\omega_{yx}\sum_{i=1}^{n}y_{i}(x_{i}^{1})^{\top}x_{t}$
$\approx y_{t}+\omega_{yx}[\sum_{i=1}^{n}y_{i}x_{i}+\omega_{xy}\sum_{i=1}^{n}y_{i}r_{i}\Sigma w^{*}]x_{t}$
$=y_{t}+\omega_{yx}(1+\omega_{xy}\sum_{i=1}^{n}r_{i}^{2})(\Sigma w^{*})^{\top}x_{t}.$
Here we used the properties of residual $r_{i}$ (in particular $\sum_{i}y_{i}x_{i}=\Sigma w^{*}$, and $\sum_{i}y_{i}r_{i}=\sum_{i}r_{i}^{2}$). Note that $(\Sigma w^{*})^{\top}x_{t}$ is basically what a gradient descent step on the original input should do. Therefore effectively, the two-layer network is doing gradient descent, but the step size is the product of $-\omega_{yx}$ and $(1+\omega_{xy}\sum_{i}r_{i}^{2})$. The factor $(1+\omega_{xy}\sum_{i}r_{i}^{2})$ depends on the level of noise, and when $\omega_{xy},\omega_{yx}<0$, the effective step size is smaller when there is more noise. This is especially helpful in the model selection problem, because intuitively one would like to perform early-stopping (small step sizes) when the noise is high.
@@ 6 Experiments
  • GD++. An even more restricted diagonal variant defined in (11).
For each experiment, we train each linear transformer modifications with a varying number of layers ($1$ to $7$) using using Adam optimizer for 200 000200000200\,000200 000 iterations with a learning rate of $0.0001$ and a batch size of 2 04820482\,0482 048. In some cases, especially for a large number of layers, we had to adjust the learning rate to prevent stability issues. We report the best result out of $5$ runs with different training seeds. We used $N=20$ in-context examples in $D=10$ dimensions. We evaluated the algorithm using 100 000100000100\,000100 000 novel sequences. All the experiments were done on a single H100 GPU with 80GB of VRAM. It took on average 4–12 hours to train a single algorithm, however experimenting with different weight decay parameters, better optimizer and learning rate schedule will likely reduce this number dramatically.
@@
Figure 2: Per-variance profile of models behavior for uniform noise variance $\sigma_{\tau}\sim U(0,\sigma_{max})$. Top two rows: 7-layer models with varying $\sigma_{max}$. Bottom row: models with varying numbers of layers, fixed $\sigma_{max}=5$. In-distribution noise is shaded gray.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: In-context learning performance for noisy linear regression across models with varying number of layers for conditional noise variance $\sigma_{\tau}\in\{1,3\}$ and $\sigma_{\tau}\in\{1,3,5\}$. Top: loss for models with various number of layers and per-variance profile for models with 7 layers. Bottom: Per-variance profile of the model across different numbers of layers. In-distribution noise is shaded gray.
@@
For the uniform noise variance, Fig. 1 shows that Full and Diag achieve comparable performance across different numbers of layers and different $\sigma_{max}$. On the other hand, GD++ converges to a higher value, closely approaching the performance of the ConstRR baseline.
@@
The top of Fig. 2 offers a detailed perspective on performance of 7-layer models and the baselines. Here, we computed per-variance profiles across noise variance range from 0 to $\sigma_{max}+1$. We can see that poor performance of GD++ comes from its inability to estimate well across the full noise variance range. Its performance closely mirrors to ConstRR, suggesting that GD++ under the hood might also be estimating a single constant variance for all the data.
AdaRR perfectly estimates problems with no noise, but struggles more as noise variance increases. TunedRR slightly improves estimation by incorporating $\sigma_{max}$ into its tunable parameters, yet its prediction suffers in the mid-range. Full and Diag demonstrate comparable performance across all noise variances. While more research is needed to definitively confirm or deny their equivalence, we believe that these models are actually not identical despite their similar performance.
At the bottom of Fig. 2 we set the noise variance to $\sigma_{max}=5$ and display a per-variance profile for models with varying layers. Two-layer models for Full and Diag behave akin to GD++, modeling only a single noise variance in the middle. However, the results quickly improve across the entire noise spectrum for 3 or more layers. In contrast, GD++ quickly converges to a suboptimal solution.
Figure 4: Weights for 4 layer linear transformer with Full parametrization trained with categorical noise $\sigma_{\tau}\in\{1,3\}$. Top: weights for $Q^{l}$ matrix, bottom: weights for $P^{l}$ matrix.
@@
Fig. 3 shows a notable difference between Diag and Full models for categorical noise variance $\sigma_{\tau}\in\{1,3\}$. This could stem from a bad local minima, or suggest a fundamental difference between the models for this problem. Interestingly, from per-variance profiling we see that Diag extrapolates better for variances not used for training, while Full, despite its lower in-distribution error, performs worse on unseen variances. Fig. 4 shows learned weights of the 4 layer linear transformer with Full parametrization. The weights are very diagonal heavy, potentially with some low-rank component.
@@
For easy comparison, we compile the results of all methods and baselines in Table 1 in the Appendix.
@@ 7 Conclusions
Our research reveals the surprising ability of linear transformers to tackle challenging in-context learning problems. We show that each layer maintains an implicit linear regression model, akin to a complex variant of preconditioned gradient descent with momentum-like behavior.
Remarkably, when trained on noisy linear regression problems with unknown noise variance, linear transformers not only outperform standard baselines but also uncover a sophisticated optimization algorithm that incorporates noise-aware step-size adjustments and rescaling. This discovery highlights the potential of linear transformers to automatically discover novel optimization algorithms when presented with the right problems, opening exciting avenues for future research, including automated algorithm discovery using transformers and generalization to other problem domains.
While our findings demonstrate the impressive capabilities of linear transformers in learning optimization algorithms, we acknowledge limitations in our work. These include the focus on simplified linear models, analysis of primarily diagonal attention matrices, and the need for further exploration into the optimality of discovered algorithms, generalization to complex function classes, scalability with larger datasets, and applicability to more complex transformer architectures. We believe these limitations present valuable directions for future research and emphasize the need for a deeper understanding of the implicit learning mechanisms within transformer architectures.
8 Acknowledgements
The authors would like to thank Nolan Miller and Andrey Zhmoginov for their valuable suggestions and feedback throughout the development of this project. Part of this work was done while Rong Ge was visiting Google Research. Rong Ge’s research is supported in part by NSF Award DMS-2031849 and CCF-1845171 (CAREER).
Appendix A Proofs from Sections 4 and  5
A.1 Proof of Theorem 4.1
@@ Appendix A Proofs from Sections 4 and  5
A.2 Proof of Lemma 4.4
This lemma is in fact a corollary of Lemma 4.3. We first give a more detailed version which explicitly state the unknown matrices $\Lambda^{l},\Gamma^{l},\Pi^{l},\Phi^{l}$:
@@
$u^{l+1}$ $=\left((1+\omega^{l}_{xy}(a^{l})^{2}\rho)I+\omega^{l}_{xx}\Sigma^{l}\right)u^{l}$
$+a^{l}\omega^{l}_{xy}\left(M^{l}+a^{l}u^{l}(w^{*})^{\top}\right)\Sigma\left(a^{l}w^{*}-w^{l}\right),$
$w^{l+1}$ $=(1+\omega^{l}_{yy}\lambda^{l})w^{l}$
$-\omega^{l}_{yx}(M^{l})^{\top}(M^{l}+a^{l}u^{l}(w^{*})^{\top})\Sigma(a^{l}w^{*}-w^{l})$
$-a^{l}\rho\omega^{l}_{yx}(M^{l})^{\top}u^{l},$
@@
$=\left(\begin{array}[]{cc}\omega^{l}_{xx}\Sigma^{l}&\omega^{l}_{xy}\alpha^{l}\\ \omega^{l}_{yx}(\alpha^{l})^{\top}&\omega^{l}_{yy}\lambda^{l}\end{array}\right).$
This implies that $A^{l}=\omega^{l}_{xx}\Sigma^{l}$, $b^{l}=\omega^{l}_{xy}\alpha^{l}$, $c^{l}=\omega^{l}_{yx}\alpha^{l}$ and $d^{l}=\omega^{l}_{yy}\lambda^{l}$. Next we rewrite $\alpha^{l}$:
@@
Now Lemma A.1 implies Lemma 4.4 immediately by setting $\Lambda^{l}=-\omega^{l}_{xy}(a^{l})^{2}\rho I-\omega^{l}_{xx}\Sigma^{l}$, $\Gamma^{l}=a^{l}\omega^{l}_{xy}\left(M^{l}+a^{l}u^{l}(w^{*})^{\top}\right)$, $s^{l}=tu{\omega}_{yy}\lambda^{l}$, $\Pi^{l}=\omega^{l}_{yx}(M^{l})^{\top}(M^{l}+a^{l}u^{l}(w^{*})^{\top})$ and $\Phi^{l}=a^{l}\rho\omega^{l}_{yx}(M^{l})^{\top}$.
A.3 Proof for Theorem 4.2
Proof.
By Theorem 4.1, we know $y^{l}_{i}=\langle w^{l},x_{i}\rangle$ for some $w^{l}$. When $n\gg d$, with high probability the norm of $y^{l}$ is on the order of $\Theta(\sqrt{n})\|w^{l}\|$, and the norm of $y^{*}$ is $\Theta(\sqrt{n})$. Therefore we only need to bound the correlation. The correlation is equal to
$|\langle y^{*},y^{l}\rangle|=|w^{l}_{1}\sum_{i=1}^{n}x_{1}^{3}+\sum_{i=1}^{n}x_{i}(1)^{2}\sum_{j=2}^{d}w^{l}_{j}x_{i}(j)|.$
We know with high probability $|\sum_{i=1}^{n}x_{i}^{3}|=O(\sqrt{n})$ because $\mathbb{E}[x_{i}^{3}]=0$. The second term can be written as $\langle w^{l},v\rangle$ where $v$ is a vector whose coordinates are $v_{1}=0$ and $v_{j}=\sum_{i=1}^{n}x_{i}(1)^{2}x_{i}(j)$ for $2\leq j\leq d$, therefore with high probability $\|v\|=O(\sqrt{nd})$. Therefore, with high probability the cosine similarity is at most
$\frac{|\langle y^{*},y^{l}\rangle|}{\|y^{*}\|\|y^{l}\|}$ $=O(1)\frac{|\langle y^{*},y^{l}\rangle|}{n\|w^{l}}$
$=O(1)\frac{|w^{l}_{1}\sum_{i=1}^{n}x_{1}^{3}+\sum_{i=1}^{n}x_{i}(1)^{2}\sum_{j=2}^{d}w^{l}_{j}x_{i}(j)|}{n\|w^{l}}$
$\leq O(1)\frac{|w^{l}_{1}||\sum_{i=1}^{n}x_{1}^{3}|+|w^{l}|\|v\|}{n\|w^{l}}$
$\leq O(1)\frac{\sqrt{n}+\sqrt{nd}}{n}.$
When $n\gg d$ this can be made smaller than any fixed constant. ∎
A.4 Proof for Theorem 5.1
In this section we prove Theorem 5.1 by finding hyperparameters for GD++ algorithm that solves least squares problems with very high accuracy. The first steps in the construction iteratively makes the data $x_{i}$’s better conditioned, and the last step is a single step of gradient descent. The proof is based on several lemma, first we observe that if the data is very well-conditioned, then one-step gradient descent solves the problem accurately:
Lemma A.2.
@@
Next we show that by applying just the preconditioning step of GD++, one can get a well-conditioned $x$ matrix very quickly. Note that the $\Sigma$ matrix is updated as $\Sigma\leftarrow(I-\gamma\Sigma)\Sigma(I-\gamma\Sigma)$, so an eigenvalue of $\lambda$ in the original $\Sigma$ matrix would become $\lambda(1-\gamma\lambda)^{2}$. The following lemma shows that this transformation is effective in shrinking the condition number
Lemma A.3.